in how many different ways can 6 friends stand together in the way Batu & Can will always be side by side?
a)60, b)120, c)180, d)240, e)360

Answer by ikleyn(52781)

.
Looking from the big distance, you can consider Batu & Can as one instance.

Then you will have 120 = 5! permutations for 5 objects.

At more detailed consideration, you can configure these two friends as Batu & Can or as Can & Batu, reversing their positions.

Then you get 240 = 2*120 = 2*5! permutations.

Answer. 240 different ways.